# KinSPEAK: Improving speech recognition for Kinyarwanda via semi-supervised learning methods

### 17 Jan 2024 v2

@@ Abstract
Despite recent availability of large transcribed Kinyarwanda speech data, achieving robust speech recognition for Kinyarwanda is still challenging. In this work, we show that using self-supervised pre-training, following a simple curriculum schedule during fine-tuning and using semi-supervised learning to leverage large unlabelled speech data significantly improve speech recognition performance for Kinyarwanda. Our approach focuses on using public domain data only. A new studio-quality speech dataset is collected from a public website, then used to train a clean baseline model. The clean baseline model is then used to rank examples from a more diverse and noisy public dataset, defining a simple curriculum training schedule. Finally, we apply semi-supervised learning to label and learn from large unlabelled data in four successive generations. Our final model achieves 3.2% word error rate (WER) on the new dataset and 15.9% WER on Mozilla Common Voice benchmark, which is state-of-the-art to the best of our knowledge. Our experiments also indicate that using syllabic rather than character-based tokenization results in better speech recognition performance for Kinyarwanda.
@@ 1 Introduction
Our approach focuses on leveraging available public domain data. First, we collect 22,000 hours of Kinyarwanda-only speech data from YouTube and use it for Self-PT pre-training. Then, we fine-tune a clean baseline model on studio-quality transcribed readings from JW.ORG website [^1] https://www.jw.org/rw/isomero/. The clean baseline model is then used to rank MCV training examples by character error rates (CER), so that we can formulate a curriculum learning schedule for a final model trained on the combined dataset. Since JW.ORG website readings are long and not segmented, we developed an easy to use mobile application that annotators used to align audio data to text, yielding about 89 hours of clean, segmented and transcribed speech data. A simple curriculum learning schedule was then devised to train an intermediate model in 6 stages, by subsequently introducing harder and harder examples to the training process. Additionally, we experiment with semi-supervised learning (Semi-SL) by transcribing and filtering audio segments from our YouTube speech data, and then adding this new dataset to our original training dataset and resuming training on the new combined dataset. This process is repeated for four generations. Our final model achieves 15.9% WER on MCV benchmark version 12 and 3.2% WER on the new dataset from JW.ORG website.
@@ 5.2 Semi-supervised learning results
Our best and final model resulted from using semi-supervised learning (Semi-SL) to leverage the unlabelled YouTube dataset. The results are presented in Table 4. We applied Semi-SL in four generations, first using top 1,000 hours of YouTube dataset, then 3,000 hours and finally 10,000 hours of data twice. This best model achieved 1.0% CER / 3.2% WER on JW.ORG test set and 4.8% CER / 15.9% WER on MCV test set.
@@
male 517 5.1 16.6 3.9 14.3
female 379 3.4 12.8 3.2 13.2
teens 182 3.2 13.3 4.7 15.4
twenties 605 3.7 13.7 3.6 13.7
thirties 135 7.1 19.4 4.3 15.5

### 02 Mar 2024 v3

@@ Abstract
Despite recent availability of large transcribed Kinyarwanda speech data, achieving robust speech recognition for Kinyarwanda is still challenging. In this work, we show that using self-supervised pre-training, following a simple curriculum schedule during fine-tuning and using semi-supervised learning to leverage large unlabelled speech data significantly improve speech recognition performance for Kinyarwanda. Our approach focuses on using public domain data only. A new studio-quality speech dataset is collected from a public website, then used to train a clean baseline model. The clean baseline model is then used to rank examples from a more diverse and noisy public dataset, defining a simple curriculum training schedule. Finally, we apply semi-supervised learning to label and learn from large unlabelled data in five successive generations. Our final model achieves 3.2% word error rate (WER) on the new dataset and 15.6% WER on Mozilla Common Voice benchmark, which is state-of-the-art to the best of our knowledge. Our experiments also indicate that using syllabic rather than character-based tokenization results in better speech recognition performance for Kinyarwanda.
@@ 1 Introduction
Our approach focuses on leveraging available public domain data. First, we collect 22,000 hours of Kinyarwanda-only speech data from YouTube and use it for Self-PT pre-training. Then, we fine-tune a clean baseline model on studio-quality transcribed readings from JW.ORG website [^1] https://www.jw.org/rw/isomero/. The clean baseline model is then used to rank MCV training examples by character error rates (CER), so that we can formulate a curriculum learning schedule for a final model trained on the combined dataset. Since JW.ORG website readings are long and not segmented, we developed an easy to use mobile application that annotators used to align audio data to text, yielding about 89 hours of clean, segmented and transcribed speech data. A simple curriculum learning schedule was then devised to train an intermediate model in 6 stages, by subsequently introducing harder and harder examples to the training process. Additionally, we experiment with semi-supervised learning (Semi-SL) by transcribing and filtering audio segments from our YouTube speech data, and then adding this new dataset to our original training dataset and resuming training on the new combined dataset. This process is repeated for five generations. Our final model achieves 15.6% WER on MCV benchmark version 12 and 3.2% WER on the new dataset from JW.ORG website.
@@ 5.2 Semi-supervised learning results
Our best and final model resulted from using semi-supervised learning (Semi-SL) to leverage the unlabelled YouTube dataset. The results are presented in Table 4. We applied Semi-SL in five generations, first using top 1,000 hours of YouTube dataset, then 3,000 hours, 10,000 hours twice and finally 20,000 hours of data. This best model achieved 1.0% CER / 3.2% WER on JW.ORG test set and 4.7% CER / 15.6% WER on MCV test set.
@@
male 517 5.1 16.5 3.8 14.0
female 379 3.4 12.7 3.2 13.2
teens 182 3.2 13.5 4.6 15.1
twenties 605 3.7 13.5 3.5 13.5
thirties 135 7.1 19.5 4.2 15.2